# sangkuriang: A pseudo-spectral Python library for Korteweg-de Vries soliton simulation

### v2

@@ 4 Conclusions
This work has presented sangkuriang, an open-source pseudo-spectral solver for the KdV equation that combines Fourier-based spatial discretization with adaptive eighth-order time integration. Validation across four test cases—ranging from isolated soliton propagation to three-body interactions—demonstrated conservation of the KdV invariants to relative errors below $10^{-5}$ for momentum and energy, while measured soliton velocities agreed with theoretical predictions to within 5%. The information-theoretic diagnostics revealed bounded spectral entropy and high phase-space determinism throughout all simulations, consistent with the integrable structure of the governing equation. Implemented in Python with Numba acceleration, the solver achieved practical throughput on modest hardware, completing the most demanding test case in under ten minutes on a laptop workstation. The combination of numerical accuracy, physical fidelity, and computational accessibility suggests that sangkuriang may serve as a useful tool for both pedagogical exploration of nonlinear wave phenomena and preliminary research investigations of soliton dynamics in dispersive media.
@@ Acknowledgments
S.H.S.H. was supported by the Dean’s Distinguished Fellowship 2023 from the University of California, Riverside. F.K., I.P.A., and D.E.I. were supported by the Research, Community Service and Innovation Program (FITB.PPMI-1-04-2025) and Early Career Research Scheme through the Directorate of Research and Innovation (2352/IT1.B07.1/TA.00/2025) from Bandung Institute of Technology.
Generative AI Disclosure
The authors declare that generative AI tools (Claude Sonnet 4.5, Anthropic) were used solely for improving English vocabulary, grammar, and readability during the preparation of this manuscript. All scientific content—including mathematical derivations, numerical implementations, software development, data analysis, interpretation of results, and figure generation–was performed entirely by the authors without AI assistance. The authors take full responsibility for the accuracy and integrity of the work presented.
Author Contributions
S.H.S.H.: Conceptualization; Formal analysis; Investigation; Methodology; Software; Validation; Visualization; Funding acquisition; Writing – original draft; Project administration. F.K.: Methodology; Funding acquisition; Writing – review & editing. I.P.A.: Funding acquisition; Supervision; Writing – review & editing. D.E.I.: Funding acquisition; Supervision; Writing – review & editing. All authors reviewed and approved the final version of the manuscript.
Open Research
The sangkuriang library source code is available on GitHub at
https://github.com/sandyherho/sangkuriang-ideal-solver under the WTFPL license and can be installed from PyPI at https://pypi.org/project/sangkuriang-ideal-solver/. The Python scripts for data analysis (https://github.com/sandyherho/suppl_sangkuriang) and all simulation outputs—including animations, NetCDF data files, figures, computational logs, and statistical analyses—archived on the Open Science Framework (OSF) (https://doi.org/10.17605/OSF.IO/MS5EJ) are released under the MIT license.

### v3

@@ 5 Conclusions
We have presented sangkuriang, an open-source pseudo-spectral solver for the KdV equation combining Fourier-based spatial discretization with adaptive eighth-order Runge–Kutta time integration. Systematic validation across four test cases of increasing complexity—from isolated soliton propagation to three-body interactions motivated by oceanic internal solitary wave dynamics—confirmed conservation of the KdV invariants to relative errors below $10^{-5}$ for momentum and energy, while measured soliton velocities agreed with the theoretical amplitude–velocity relation to within 5%. Information-theoretic diagnostics revealed bounded spectral entropy and near-unit phase-space determinism throughout all simulations, providing quantitative evidence that the numerical solutions preserve the integrable Hamiltonian structure of the governing equation. Implemented in Python with Numba-accelerated parallel execution, the solver completed the most demanding configuration—a three-soliton interaction over 80 s at $N=1024$ resolution—in under ten minutes on a laptop workstation, demonstrating accessibility for researchers in oceanography and GFD without requiring high-performance computing resources.
The idealized test cases examined here—single-soliton propagation, symmetric pairs, overtaking collisions, and rank-ordered three-wave packets—capture the essential nonlinear dispersive dynamics that underpin a wide range of oceanic solitary wave phenomena, from internal soliton trains on continental shelves to long gravity waves in coastal basins. By establishing the fidelity of the sangkuriang solver against these canonical benchmarks, the present work provides a verified computational foundation upon which more realistic ocean wave studies can be built.
Beyond its immediate utility for KdV simulations, the diagnostic framework developed here—combining conservation monitoring, spectral information measures, and recurrence quantification—offers a transferable methodology for assessing integrability preservation in numerical solutions of nonlinear dispersive PDEs arising in GFD. Extensions to variable-coefficient KdV variants relevant to wave propagation over realistic bathymetry, higher-order dispersive equations such as the Kawahara equation, rotational generalizations including the Ostrovsky equation for internal waves influenced by the Earth’s rotation, and two-dimensional formulations such as the Kadomtsev–Petviashvili equation represent natural directions for future development within the existing pseudo-spectral framework. Such extensions would progressively bridge the gap between the idealized dynamics studied here and the full complexity of nonlinear wave propagation in the real ocean.
Author Contributions
Conceptualization, D.E.I. and S.H.S.H.; methodology, D.E.I., S.H.S.H., and F.K.; software, S.H.S.H.; validation, D.E.I., S.H.S.H., and F.K.; formal analysis, S.H.S.H.; investigation, D.E.I. and S.H.S.H.; resources, D.E.I., I.P.A., and R.S.; data curation, S.H.S.H.; writing—original draft preparation, S.H.S.H.; writing—review and editing, D.E.I., A.P., R.D.K., F.K., I.P.A., K.A.S., A.P.H., F.R.F., and R.S.; visualization, S.H.S.H.; supervision, D.E.I., I.P.A., K.A.S., A.P.H., F.R.F., and R.S.; project administration, D.E.I.; funding acquisition, D.E.I., S.H.S.H., F.K., A.P.H., F.R.F., and R.S. All authors have read and agreed to the published version of the manuscript.
Funding
This study was supported by the Program Penelitian, Pengabdian kepada Masyarakat, dan Inovasi FITB 2026, the Dean’s Distinguished Fellowship 2023 from the University of California, Riverside, the Riset Dosen Muda 2025, the Riset Unggulan 2026, and the Riset Talenta Unggul 3P 2026 from Bandung Institute of Technology.
Data Availability
The sangkuriang library source code is available on GitHub at
https://github.com/sandyherho/sangkuriang-ideal-solver under the MIT license and can be installed from PyPI at https://pypi.org/project/sangkuriang-ideal-solver/. The Python scripts for data analysis (https://github.com/sandyherho/suppl_sangkuriang) and all simulation outputs—including animations, NetCDF data files, figures, computational logs, and statistical analyses—archived on the Open Science Framework (OSF) (https://doi.org/10.17605/OSF.IO/MS5EJ) are released under the MIT license.
@@ Acknowledgments
The authors thank the anonymous reviewers for their constructive comments. During the preparation of this work, the authors used Claude Sonnet 4.5 (Anthropic) to assist with English grammar, vocabulary, and stylistic refinement. All mathematical derivations, numerical methods, scientific analysis, interpretation of results, and intellectual content were produced entirely by the authors. The authors reviewed and edited all AI-assisted outputs and take full responsibility for the content of the publication.
Conflicts of Interest
The authors declare no conflicts of interest. The funders had no role in the design of the study; in the collection, analyses, or interpretation of data; in the writing of the manuscript; or in the decision to publish the results.
Abbreviations
CF Climate and Forecast
CPU Central Processing Unit
DFT Discrete Fourier Transform
DOP853 Dormand–Prince 8(5,3) method
FFT Fast Fourier Transform
FFTW Fastest Fourier Transform in the West
GFD Geophysical Fluid Dynamics
IFFT Inverse Fast Fourier Transform
ISW Internal Solitary Wave
JIT Just-In-Time
KdV Korteweg–de Vries
LLVM Low Level Virtual Machine
LMC López–Ruiz–Mancini–Calbet
MKL Math Kernel Library
NetCDF Network Common Data Format
ODE Ordinary Differential Equation
OpenMP Open Multi-Processing
PDE Partial Differential Equation
PyPI Python Package Index
RMS Root Mean Square
RQA Recurrence Quantification Analysis